in a college class of 250 graduating seniors,
50 have job waiting 10 are going to medical school,
20 are gong to law school, and
80 are going to various other kinds of graduate schools.
select one graduate at random
a)what is the probability that the student is going to graduate school?
b)what is the probability that the student is going to medical school?
c)what is the probability that the student does not continue in school?

Answer by Edwin McCravy(20062)

in a college class of 250 graduating seniors,
50 have job waiting
10 are going to medical school,
20 are gong to law school,
and 80 are going to various other kinds of graduate schools.
select one graduate at random
a)what is the probability that the student is going to graduate school?
80 out of 250 is 80/250 or 8/25

b)what is the probability that the student is going to medical school?
10 out of 250 is 1/25

c)what is the probability that the student does not continue in school?
10+20+80 = 110 are continuing in school

250-110 = 140 are not continuing in school

140 out of 250 = 140/250 = 14/25.

Edwin